OpenStudy (anonymous):

n(5n - 2)(2n + 5) = 0

OpenStudy (dumbcow):

solving for n

OpenStudy (dumbcow):

n=0 5n-2=0 5n = 2 n = 2/5 2n+5=0 2n = -5 n =-5/2